OpenStudy (anonymous):

find the derivative of ((4x-2rootx)/2) using power, product, quotient rules

OpenStudy (amistre64):

(1/2)(4x -2(x^1/2)) (1/2)(4 - x^(-1/2) 2 - 1/2sqrt(x)

OpenStudy (amistre64):

maybe

OpenStudy (amistre64):

\[\frac{4x}{2} - \frac{2\sqrt{x}}{2}\]

OpenStudy (amistre64):

Dx (2x - sqrt(x)) = 2 - 1/2sqrt(2) yep